OpenStudy (idealist10):

Find an integrating factor of the form u(x, y)=P(x)Q(y) for (a*cosxy-y*sinxy)dx+(b*cosxy-x*sinxy)dy=0 and solve the equation.

OpenStudy (idealist10):

\[F(x)(b \cos xy-x \sin xy)-G(y)(a \cos xy-y \sin xy)\]

OpenStudy (idealist10):

What's the partial derivative of (a*cosxy-y*sinxy) with respect to y?

OpenStudy (idealist10):

Is it -sinxy?

OpenStudy (anonymous):

\[\begin{align*} F(x)(b\cos xy-x\sin xy)\quad&\\ -G(y)(a\cos xy-y\sin xy)&=(a\cos xy-y\sin xy)_y-(b\cos xy-x\sin xy)_x \end{align*}\] \[\frac{\partial}{\partial y}(a\cos xy-y\sin xy)=-ax\sin(xy)-\sin xy-xy\cos xy\\ \frac{\partial }{\partial x}(b\cos xy-x\sin xy)=-by\sin xy-\sin xy-xy\cos xy\]

OpenStudy (anonymous):

Sorry about the random set of parentheses around the \(xy\)...

OpenStudy (idealist10):

That's fine.

OpenStudy (anonymous):

Do those derivatives make sense? It's mostly chain and product rules.

OpenStudy (idealist10):

Yes, they do make sense.

OpenStudy (idealist10):

I have a question. What's the prerequisite of Complex Variables?

OpenStudy (anonymous):

Hmm, that depends... If it's some sort of rigorous course on complex analysis, you'll need to be familiar with a bit of real analysis and a lot of calculus. If it's to do with differential equations, it's mostly a matter of knowing Euler's formula, so trigonometry and a bit of linear algebra. Or if you mean complex variables in general, it's essentially an extension of algebra and deals with equations in two variables.

OpenStudy (idealist10):

So after learning Partial Differential Equations, should I learn Complex Variables?

OpenStudy (anonymous):

Anyway... \[\begin{align*} F(x)(b\cos xy-x\sin xy)\quad&\\ -G(y)(a\cos xy-y\sin xy)&=(-ax\sin xy-\sin xy-xy\cos xy)\\&\quad-(-by\sin xy-\sin xy-xy\cos xy)\\\\ F(x)(b\cos xy-x\sin xy)\quad&\\ -G(y)(a\cos xy-y\sin xy)&=(by-ax)\sin xy\\\\ F(x)&=\frac{(by-ax)\sin xy+G(y)(a\cos xy-y\sin xy)}{b\cos xy-x\sin xy} \end{align*}\] Hmm... might take a while to figure out this "constant" function we need to eliminate the appropriate variables...

OpenStudy (idealist10):

How did you get (by-ax)sinxy?

OpenStudy (anonymous):

You subtract the partial derivatives. The \(-\sin xy\) and \(-xy\cos xy\) terms disappear, and I factored the common \(\sin xy\).

OpenStudy (anonymous):

I have to get going, but I'll get back to this when I have the time.

OpenStudy (idealist10):

So what should the "constant" function be?

OpenStudy (anonymous):

Finally got it after some effort. The integrating factor, I mean, not the \(C\) function. I don't think that approach will work here. Instead, what I did was assume \[\large\begin{cases} F(x)=\alpha x^\gamma\\\\ G(y)=\beta y^\delta \end{cases}\] which leads to getting the IF, \[\large u(x,y)=e^{ax+by}\]

OpenStudy (anonymous):

I'd post the details now, but there's a lot of writing involved... I'll return to this another time if you need an in-depth explanation.

OpenStudy (idealist10):

How did you get F(x) and G(y)? Can you please show your work?

OpenStudy (anonymous):

Guesswork, mostly... I started off assuming \(\large F(x)=\alpha x^\gamma\) and \(\large G(y)=\beta y^\delta\) and tried to see if I could find constants \(\alpha,\beta,\gamma,\delta\) that would work.

OpenStudy (anonymous):

Substituting those "solutions" into the equation: \[\large\begin{align*} \alpha x^\gamma(b\cos xy-x\sin xy)\quad&\\ -\beta y^\delta(a\cos xy-y\sin xy)&=(by-ax)\sin xy\\\\ \alpha bx^\gamma\cos xy-\alpha x^{\gamma+1}\sin xy\quad&\\ -\beta ay^\delta\cos xy+\beta y^{\delta+1}\sin xy&=(by-ax)\sin xy\\\\ (\alpha bx^\gamma-\beta ay^\delta)\cos xy\quad\quad\quad&\\+(\beta y^{\delta+1}-\alpha x^{\gamma+1})\sin xy&=(by-ax)\sin xy \end{align*}\] Matching up sines and cosine gives \[\large\begin{cases} \alpha bx^\gamma-\beta ay^\delta=0\\ \beta y^{\delta+1}-\alpha x^{\gamma+1}=by-ax \end{cases}\] The first equation needs to have \(\gamma=\delta=0\) so that we can remove the \(x\) and \(y\). Solving for \(\alpha\) then gives \[\alpha=\frac{\beta a}{b}\] and subbing into the second equation gives \[\beta y-\frac{\beta a}{b} x=by-ax\] which gives means \(\beta=b\) and \(\dfrac{\beta a}{b}=\alpha~~\iff~~\alpha=a\). Hence \[\large\begin{cases} F(x)=\alpha\\\\ G(y)=\beta \end{cases}\] and the IF becoems \[\large\log u(x,y)=\int \alpha ~dx+\int\beta~dy~~\iff~~u(x,y)=e^{ax+by}\]

OpenStudy (anonymous):

Oh and I never replied to your question about your complex-number course. I wouldn't know how to advise you on what courses to take. It all really depends on what you want to do for a living. I've never taken a course on PDEs, but I have had a semester of complex analysis, and it had nothing to do with differential equations. Like I mentioned before, it was all calculus, but I couldn't quite see much practical application to it.

OpenStudy (anonymous):

There's actually another way of solving this without finding the integrating factor. I know the instructions require finding the IF, but I deemed it worth mentioning: \[\begin{align*}(a \cos xy-y \sin xy)+(b \cos xy-x \sin xy)y'&=0\\\\ (a+by')\cos xy&=(y+xy')\sin xy\\\\ (a+by')\cos xy&=\frac{d}{dx}[-\cos xy]\\\\ a+by'&=-\frac{\dfrac{d}{dx}[\cos xy]}{\cos xy}\\\\ a+by'+\frac{\dfrac{d}{dx}[\cos xy]}{\cos xy}&=0\\\\ \int\left(a+by'+\frac{\dfrac{d}{dx}[\cos xy]}{\cos xy}\right)~dx&=C\\\\ ax+by+\ln(\cos xy)&=C\end{align*}\]

OpenStudy (idealist10):

Thank you so much for the extra info and the guide on math courses.